OpenStudy (anonymous):

Hey really need help on this Geometry question! please any one out there to help ASAP?

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

so i need to find the Total Area of this .. i know the formula which is T.A= ph+2B but when i did that my answer was not one of the options

OpenStudy (anonymous):

what grade are you in I am in 8th

OpenStudy (anonymous):

im a sophomore

OpenStudy (anonymous):

oh sorry I can't help you

OpenStudy (anonymous):

do you know any one who could?

OpenStudy (anonymous):

I can tag @Directrix

OpenStudy (anonymous):

would you

OpenStudy (anonymous):

I just did

OpenStudy (anonymous):

thanks

Directrix (directrix):

Is this a cube?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I had done TA= ph + 2B

Directrix (directrix):

The total surface area is the area of one square face times 6. So, 12 times 12 times 6 = what

OpenStudy (anonymous):

864

OpenStudy (anonymous):

but where did you get the 6 from?

Directrix (directrix):

A cube has 6 faces.

Directrix (directrix):

This should work: TA= ph + 2B TA = 48 * 12 + 2 * 12 * 12 =

OpenStudy (anonymous):

so would 864 be the answer then?

Directrix (directrix):

Your formula barfed up 864

OpenStudy (anonymous):

that = 604

Directrix (directrix):

Where did you make the area on your formula.

OpenStudy (anonymous):

in the formula they gave me it was Total Area= ph + 2B and P= to the perimeter of the bases, oohhhh i think i see where i went wrong in my work.. i multiplied not adding

OpenStudy (anonymous):

ok i got it thanks for your help :)

Directrix (directrix):

You are welcome.